# Possible detection of HFQPOs associated with ‘unknown’ variability class of GRS 1915+105

### v1

@@ 1 Introduction
Table 1: Observation details of GRS $1915+105$ observed by AstroSat and NuSTAR during June, $2017$. Column 11, 22, 44, 55 and $6$ are for ObsID, Epoch, Orbit number, MJD and exposure time. In column 33, ‘Seg.’ denotes the light curve segments for Epochs AS11 and AS22. The mean detected ($r_{\rm det}$) and incident ($r_{\rm in}$) count rates, along with the mean hardness ratios (HRs) over the entire exposures for each variability class, are tabulated. $\rm r_{det}$_var, HR1_var and HR2_var are the variance of the respective quantities normalized with the corresponding average values. In column 111211-12, the variability classes and the presence of Low//High-frequency QPO features in respective observations are given. See the text for details.
@@ 2 Observation and Data Reduction
In the present work, we analyze two Guaranteed Time (GT) observations of GRS 1915++105 during June $2017$ with SXT and LAXPC onboard AstroSat Agrawal et al. (2017). Three more adjacent observations available with AstroSat and NuSTAR during May-July, $2017$ are also analyzed. Additionally, we analyze three RXTE/PCA observations (RX00-RX22) in our study, which are available in the HEASARC archive [^1] https://heasarc.gsfc.nasa.gov/db-perl/W3Browse/w3browse.pl. The details of all these observations are tabulated in Table 1 along with Epoch numbers. The Epochs of these observations are marked using vertical lines in the MAXI/GSC light curve (Fig. 1).
We extract LAXPC level-1 data in event analysis mode, available in AstroSat public archive [^2] https://webapps.issdc.gov.in/astro_archive/archive/Home.jsp using the standard data reduction software LaxpcSoftv3.4.4333http://www.tifr.res.in/~astrosat_laxpc/LaxpcSoft.html. The details of LAXPC data extraction procedure and analysis methods are mentioned in Sreehari et al. (2019, 2020); Majumder et al. (2022). It is worth mentioning that, due to the constraints of Earth occultation and passage of the satellite in the region of South Atlantic Anomaly (SAA), continuous observation is not possible with AstroSat. Subsequently, the data corresponding to a given observation of AstroSat/LAXPC are generally divided into multiple Good Time Interval (GTI) segments of the observation time using the analysis software. For example, the entire data of AS1 observation contains $4$ GTI segments ($1a$, $1b$, $1c$ and $1d$) (see Table 1), which are distributed over $4$ orbits of observation. Further, we mention that the timing analyses including the study of variability properties and power density spectra are carried out for such individual segments excluding the data gaps, whereas the spectral analysis is performed considering the entire observation data of AS1 and AS2.
We utilize both LAXPC10 and LAXPC20 data for the timing analyses, whereas only LAXPC20 data are used in the spectral study following Antia et al. (2021); Majumder et al. (2022). Note that, the data from the top layer of the detectors and that corresponds to the single events are considered for the analysis. Accordingly, we generate the source and background spectra along with instrument response files for LAXPC20 following Antia et al. (2017). We analyze the SXT data following the guidelines provided by the SXT instrument team [^4] https://www.tifr.res.in/~astrosat_sxt/index.html. The SXT light curves, spectra and ancillary response files are extracted using XSELECT V2.5b from a $12$ arcmin circular source region using the level-2 cleaned event files available at ISSDC [^5] https://webapps.issdc.gov.in/astro_archive/archive/Home.jsp. data archive.
@@ 3.1 MAXI/GSC Monitoring of GRS 1915+105
We study the evolution of the source GRS $1915+105$ with daily MAXI/GSC monitoring. In Fig. 1a, we present the light curve in 2202-20 keV energy range during MJD 577115806157711-58061. In Fig. 1b, we show the variation of hardness ratio (HR) defined as the ratio of counts in 6206-20 keV and 262-6 keV energy bands. The dashed vertical lines in brown, black, green and magenta denote the Epochs corresponding to AstroSat and NuSTAR observations under consideration. Further, we carry out a correlation study between HR and count rates for two distinct regions of the MAXI/GSC light curves, represented by orange (Seg 1) and blue (Seg 2) colors in Fig. 1. The variation of HR against the count rates for each of these two segments is shown in the insets of Fig. 1b. While computing the Pearson correlation coefficient ($p$) of these distributions, we consider all the data points in Seg 1, whereas count rates exceeding $5$ counts $\rm cm^{-2}$ $\rm s^{-1}$ are excluded for Seg 2. With this, we find a moderate negative (positive) correlation between HR and count rate for Seg 1 (Seg 2), with p=0.7(+0.7)p=-0.7(+0.7).
@@ 5.1 An ‘Unknown’ Variability Class in GRS 1915+105
Meanwhile, several studies are carried out to explain the origin of $\rho$ class variability in GRS $1915+105$, relying on the classical disc instability model Lightman & Eardley (1974) for a truncated accretion disc and compact corona configuration Honma et al. (1991); Taam et al. (1997); Szuszkiewicz & Miller (1998); Janiuk et al. (2000). Further, Massa et al. (2013) argued that the limit cycle behaviour, which mostly depends on the accretion rate, is typical for the $\rho$ class variability and explains the hard lags as well. Moreover, Nandi et al. (2000) proposed that the bursts in the light curve of $\rho$ class are closely associated with the so-called ‘On’ state of the source. Most importantly, they found the features of the small-scaled version of a structured $\rho$ class in the light curve of canonical $\kappa$ class variability. This essentially indicates the intrinsic connection between these variability signatures, perhaps produced from the interplay of inner Comptonizing corona and the feedback of disc winds Nandi et al. (2000); Chakrabarti & Nandi (2000). More recently, a phenomenological non-linear mathematical model has also been put forward to describe the complex variability classes of GRS $1915+105$ Massaro et al. (2020). However, we find that the model is unable to reproduce the structured variability signatures consisting of irregular low-count dips and spike-like features as observed in the light curve of the ‘unknown’ variability class. These findings further indicate that the ‘unknown’ variability class ($\tau$) is quite different from the canonical $\chi$ and $\rho$ class and possibly could not be explained as a transition between these two classes.

### v2

@@ 1 Introduction
Table 1: Observation details of GRS $1915+105$ observed by AstroSat and NuSTAR during June, $2017$. Column 11, 22, 44, 55 and $6$ are for ObsID, Epoch, Orbit number, MJD and exposure time. In column 33, ‘Seg.’ denotes the light curve segments for Epochs AS11 and AS22. The mean detected ($r_{\rm det}$) and incident ($r_{\rm in}$) count rates, along with the mean hardness ratios (HRs) over the entire exposures for each variability class, are tabulated. HR1_var, HR2_var and $\rm r_{det}$_var are the fractional variabilities of the respective quantities. In column 141514-15, the variability classes and the presence of Low//High-frequency QPO features in respective observations are given. See the text for details.
@@ 2 Observation and Data Reduction
In the present work, we analyze two Guaranteed Time (GT) observations of GRS 1915++105 during June $2017$ with SXT and LAXPC onboard AstroSat Agrawal (2006); Singh et al. (2014); Agrawal et al. (2017). Three more adjacent observations available with AstroSat and NuSTAR during May-July, $2017$ are also analyzed. Additionally, we analyze three RXTE/PCA observations (RX00-RX22) in our study, which are available in the HEASARC archive [^1] https://heasarc.gsfc.nasa.gov/db-perl/W3Browse/w3browse.pl. The details of all these observations are tabulated in Table 1 along with Epoch numbers. The Epochs of these observations are marked using vertical lines in the MAXI/GSC light curve (Fig. 1).
We extract LAXPC Antia et al. (2017) level-1 data in event analysis mode, available in AstroSat public archive [^2] https://webapps.issdc.gov.in/astro_archive/archive/Home.jsp using the standard data reduction software LaxpcSoftv3.4.4333http://www.tifr.res.in/~astrosat_laxpc/LaxpcSoft.html. The details of LAXPC data extraction procedure and analysis methods are mentioned in Sreehari et al. (2019, 2020); Majumder et al. (2022). It is worth mentioning that, due to the constraints of Earth occultation and passage of the satellite in the region of South Atlantic Anomaly (SAA), continuous observation is not possible with AstroSat. Subsequently, the data corresponding to a given observation of AstroSat/LAXPC are generally divided into multiple Good Time Interval (GTI) segments of the observation time using the analysis software. For example, the entire data of AS1 observation contains $4$ GTI segments ($1a$, $1b$, $1c$ and $1d$) (see Table 1), which are distributed over $4$ orbits of observation. Further, we mention that the timing analyses including the study of variability properties and power density spectra are carried out for such individual segments excluding the data gaps, whereas the spectral analysis is performed considering the entire observation data of AS1 and AS2.
We utilize both LAXPC10 and LAXPC20 data for the timing analyses, whereas only LAXPC20 data are used in the spectral study following Antia et al. (2021); Majumder et al. (2022). Note that, the data from the top layer of the detectors and that corresponds to the single events are considered for the analysis. Accordingly, we generate the source and background spectra along with instrument response files for LAXPC20 following Antia et al. (2017). We analyze the SXT Singh et al. (2016, 2017) data following the guidelines provided by the SXT instrument team [^4] https://www.tifr.res.in/~astrosat_sxt/index.html. The SXT light curves, spectra and ancillary response files are extracted using XSELECT V2.5b from a $12$ arcmin circular source region using the level-2 cleaned event files available at ISSDC [^5] https://webapps.issdc.gov.in/astro_archive/archive/Home.jsp. data archive.
@@ 3.1 MAXI/GSC Monitoring of GRS 1915+105
We study the evolution of the source GRS $1915+105$ with daily MAXI/GSC monitoring. In Fig. 1a, we present the light curve in 2202-20 keV energy range during MJD 577115806157711-58061. In Fig. 1b, we show the variation of hardness ratio (HR) defined as the ratio of counts in 6206-20 keV and 262-6 keV energy bands. The dashed vertical lines in brown, black, green and magenta denote the Epochs corresponding to AstroSat and NuSTAR observations under consideration. Further, we carry out a correlation study between HR and count rates for two distinct regions of the MAXI/GSC light curves, represented by orange (Seg 1) and blue (Seg 2) colors in Fig. 1. The variation of HR against the count rates for each of these two segments is shown in the insets of Fig. 1b. While computing the Pearson correlation coefficient ($p$) of these distributions, we consider all the data points in Seg 1, whereas count rates exceeding $5$ counts $\rm cm^{-2}$ $\rm s^{-1}$ are excluded for Seg 2. With this, we find a strong negative (positive) correlation between HR and count rate for Seg 1 (Seg 2) with p=0.7(+0.7)p=-0.7\;(+0.7) for one-tailed chance probability of $3.1\times 10^{-16}$ ($5.1\times 10^{-15}$).
@@ 5.1 An ‘Unknown’ Variability Class in GRS 1915+105
Meanwhile, several studies are carried out to explain the origin of $\rho$ class variability in GRS $1915+105$, relying on the classical disc instability model Lightman & Eardley (1974) for a truncated accretion disc and compact corona configuration Honma et al. (1991); Taam et al. (1997); Szuszkiewicz & Miller (1998); Janiuk et al. (2000). Further, Massa et al. (2013) argued that the limit cycle behaviour, which mostly depends on the accretion rate, is typical for the $\rho$ class variability and explains the hard lags as well. Moreover, Nandi et al. (2000) proposed that the bursts in the light curve of $\rho$ class are closely associated with the so-called ‘On’ state of the source. Most importantly, they found the features of the small-scaled version of a structured $\rho$ class in the light curve of canonical $\kappa$ class variability. This essentially indicates the intrinsic connection between these variability signatures, perhaps produced from the interplay of inner Comptonizing corona and the feedback of disc winds Nandi et al. (2000); Chakrabarti & Nandi (2000). A phenomenological non-linear mathematical model has also been put forward to describe the complex variability classes of GRS $1915+105$ Massaro et al. (2020a, b). This model contains an input function $J(t)$ which is used to reproduce light curves of various classes. For example, constant input values Massaro et al. (2020a) reproduce the classes $\chi$, $\delta$ and $\rho$, while several other classes ($\alpha$, $\gamma$, $\lambda$, $\kappa$, $\omega$, $\xi$, and $\theta$) result when rather simple step function or sawtooth modulations of $J(t)$ are considered Massaro et al. (2020b). We find that the model with a constant input function is unable to reproduce the structured variability signatures consisting of irregular low-count dips and spike-like features as observed in the light curve of the ‘unknown’ variability class ($\tau$). These findings further suggest that the class $\tau$ is quite different from the canonical $\chi$ and $\rho$ classes and possibly could not be explained as a transition between these two classes. However, we cannot exclude the possibility of getting a functional form of $J(t)$, which could model the $\tau$ class variability. It is noteworthy that Alberti et al. (2022) also proposed a formalism based on a stochastic dynamical process to describe the complex variability signatures of GRS $1915+105$. However, applying such models to the present data is beyond the scope of this work and will be addressed elsewhere.